# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

### v1

@@ 2.2 Mini-batch NSHB and mini-batch SHB
Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximated parameter of DNN. Then, mini-batch NSHB uses $b_{t}$ loss functions $f_{\xi_{t,1}},\cdots,f_{\xi_{t,b_{t}}}$ randomly chosen from $\{f_{1},\cdots,f_{n}\}$ at each step $t$, where $\bm{\xi}_{t}=(\xi_{t,1},\cdots,\xi_{t,b_{t}})^{\top}$ is independent of $\bm{\theta}_{t}$ and $b_{t}$ is a batch size satisfying $b_{t}\leq n$. The Mini-batch NSHB optimizer is as in Algorithm 1.
@@
0:  $\bm{\theta}_{0},\bm{m}_{-1}:=\bm{0}$ (initial point), $b_{t}>0$ (batch size), $\eta>0$ (learning rate), $\beta\in[0,1)$ (momentum weight), $T\geq 1$ (steps)
0:  $(\bm{\theta}_{t})\subset\mathbb{R}^{d}$
@@
where $\beta\in[0,1)$ and $\alpha>0$. SHB defined by (2) with $\beta=0$ coincides with SGD. SHB defined by (2) has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Meanwhile, Algorithm 1 is called the normalized-SHB (NSHB) optimizer (Gupal & Bazhenov, 1972) and has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Hence, NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2).
@@ 3.1 Constant batch size scheduler
[…]
does not always converge to 00 (a proof of Theorem 3.1 is given in Appendix A.3).
@@
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that
$\eta\leq\max\left\{\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{2}}{(5-\beta+2\beta^{2})L}\right\},$
and Constant BS defined by (3), where $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$ and $f^{\star}$ is the minimum value of $f$ over $\mathbb{R}^{d}$ (see (A1)). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{L\eta\sigma^{2}}{b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
@@
From the discussion in Section 2.2 such that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2), Theorem 3.1 leads to the following.
Theorem 3.2 (Upper bound of $\min_{t}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|$ of mini-batch SHB with Constant BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\alpha\leq\max\left\{\frac{(1-\beta)^{2}}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{3}}{(5-\beta+2\beta^{2})L}\right\},$
and Constant BS defined by (3). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}+\frac{L\alpha\sigma^{2}}{(1-\beta)b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
that is,
@@ 3.2 Increasing batch size scheduler
An example of $b_{t}$ (Smith et al., 2018; Umeda & Iiduka, 2024) is, for all $m\in[0:M]$ and all $t\in S_{m}=\mathbb{N}\cap[\sum_{k=0}^{m-1}K_{k}E_{k},\sum_{k=0}^{m}K_{k}E_{k})$ ($S_{0}:=\mathbb{N}\cap[0,K_{0}E_{0})$),
[…]
where $\delta>1$, and $E_{m}$ and $K_{m}$ are the numbers of, respectively, epochs and steps per epoch when the batch size is $\delta^{m}b_{0}$. For example, the exponential growth batch size defined by (4) with $\delta=2$ makes batch size double each $E_{m}$ epochs. We may modify the parameters $a$ and $\delta$ to $a_{t}$ and $\delta_{t}$ monotone increasing with $t$. The total number of steps for the batch size to increase $M$ times is $T=\sum_{m=0}^{M}K_{m}E_{m}$.
@@
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that
@@
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}$
$+\frac{2L\eta\sigma^{2}K_{\max}E_{\max}\delta}{(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
@@
From the discussion in Section 2.2 such that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2), Theorem 3.3 leads to the following convergence rate of SHB defined by (2) with an increasing batch size.
Theorem 3.4 (Convergence of mini-batch SHB with Exponential Growth BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}$
$+\frac{2L\alpha\sigma^{2}K_{\max}E_{\max}\delta}{(1-\beta)(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
that is,
Here, we sketch a proof of Theorem 3.3 (a detailed proof of Theorem 3.3 is given in Appendix A.2).
@@
    We first show that an upper bound of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ is $\frac{\sigma^{2}}{b_{t}}$ (Proposition A.1) and then that an upper bound of the variance of $\bm{m}_{t}=(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f_{B_{i}}(\bm{\theta}_{i})$ is $(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ (Lemma A.2) using the idea of the proof of (Liu et al., 2020, Lemma 1).
@@
    We next show that an auxiliary point $\bm{z}_{t}=\frac{1}{1-\beta}\bm{\theta}_{t}-\frac{\beta}{1-\beta}\bm{\theta}_{t-1}$ ($t\geq 1$), which is used to analyze SGDM (Yan et al., 2018; Yu et al., 2019; Liu et al., 2020), satisfies $\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]\leq f(\boldsymbol{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}$ using the descent lemma (see (7)). Using the Cauchy–Schwarz inequality, Young’s inequality, and the upper bound of the variance of $\bm{m}_{t}$ (Lemma A.2) provides an upper bound of $-\eta\mathbb{E}[X_{t}]$ (see (12)). An upper bound of $\mathbb{E}[Y_{t}]$ is provided by using the upper bound $\frac{\sigma^{2}}{b_{t}}$ of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ (see Lemma A.3 for details of the upper bounds of $-\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$).
@@
    We define the Lyapunov function $L_{t}$ by $L_{t}=f(\bm{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\|^{2}$, where $c_{i}$ is defined as in Lemma A.4. Using the above upper bounds of $-\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$, we have that $\mathbb{E}[L_{t+1}-L_{t}]\leq-D\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+U_{t}$ (Lemma A.4), where $D\in\mathbb{R}$ depends on $\eta$, $\beta$, and $c_{1}$, and $U_{t}>0$ depends on $\sigma^{2}$, $b_{t}$, and $c_{1}$.
@@
    Setting $\eta$ to satisfy (5) leads to the finding that $D\geq\frac{\eta}{2}>0$ and $U_{t}\leq L\eta^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ (see (19) and (20)). As a result, we have that
@@ 3.3 Setting of hyperparameter $\delta$ in Exponential Growth BS (4)
Let $\eta$ and $\beta$ be fixed in Algorithm 1 (e.g., $\eta=0.1$ and $\beta=0.9$). Then, Theorems 3.3 and 3.4 indicate that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for each of mini-batch NSHB and mini-batch SHB with Exponential Growth BS (4) is $O(\sqrt{\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}+\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}})$, which implies that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ is (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). In Section 4, we verify whether this theoretical result holds in practice.
3.4 Comparisons of our convergence results with previous ones
@@
Let us compare Theorems 3.13.4 with the previous results listed in Table 1. Theorem 1 in (Liu et al., 2020) ((6) in Table 1) indicated that NSHB using a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$. Since the upper bound $O(\sqrt{\frac{1}{T}+\sigma^{2}})$ converges to $O(\sigma)>0$ when $T\to+\infty$, NSHB in this case does not always converge to stationary points of $f$. The result in (Liu et al., 2020) coincides with Theorem 3.1 indicating that NSHB has $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$ in the sense that NSHB using a constant learning rate and momentum weight does not converge to stationary points of $f$111(Liu et al., 2020, Theorem 3) was a convergence of multistage SGDM. However, since the proof of (Liu et al., 2020, (60), Pages 35 and 36) might not hold for $\beta_{i}<1$, the theorem does not apply here.. Corollary 1 in (Yu et al., 2019) ((5) in Table 1) indicated that SHB using constant $\eta=O(\frac{1}{\sqrt{T}})$ and constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/4}})$. Using $\eta=O(\frac{1}{\sqrt{T}})$ is necessary to set the number of steps $T$ before implementing SHB. Since the $T$ is fixed, we cannot diverge $T$, that is, the upper bound $O(\frac{1}{T^{1/4}})$ for SHB is a fixed positive constant and does not converge to 00. Meanwhile, Theorem 3.2 is the result for SHB using a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight, which can be obtained by Theorem 3.1. Theorem 3.2 indicates that SHB has $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$. Hence, Theorem 3.2 coincides with the result in Corollary 1 in (Yu et al., 2019) in the sense that SHB with constant learning rate does not always converge to stationary points of $f$.
@@
Theorems 1 and 2 in (Gitman et al., 2019) ((2) and (3) in Table 1) indicated that QHM, which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ or an increasing momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$. Our results in Theorems 3.3 and 3.4 guarantee the convergence of NSHB and SHB with constant learning rate $\eta=O(\frac{1}{L})$, constant momentum weight $\beta$, and an increasing batch size $b_{t}$ in the sense of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{\sqrt{T}})$.

### v2

@@ 2.2 Mini-batch NSHB and mini-batch SHB
Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximated parameter of DNN. Then, mini-batch NSHB uses $b_{t}$ loss functions $f_{\xi_{t,1}},\cdots,f_{\xi_{t,b_{t}}}$ randomly chosen from $\{f_{1},\cdots,f_{n}\}$ at each step $t$, where $\bm{\xi}_{t}=(\xi_{t,1},\cdots,\xi_{t,b_{t}})^{\top}$ is independent of $\bm{\theta}_{t}$ and $b_{t}$ is a batch size satisfying $b_{t}\leq n$. The Mini-batch NSHB optimizer is listed in Algorithm 1.
where $\beta\in[0,1)$ and $\alpha>0$. SHB defined by (2) with $\beta=0$ coincides with SGD. SHB defined by (2) has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Meanwhile, Algorithm 1 is called the normalized-SHB (NSHB) optimizer (Gupal and Bazhenov, 1972) and has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Hence, NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2).
@@
$𝟎\bm{\theta}_{0},\bm{m}_{-1}:=\bm${0} (initial point), $b_{t}>0$ (batch size), $\eta>0$ (learning rate), $\beta\in[0,1)$ (momentum weight), $T\geq 1$ (steps)
[…]
@@ 3.1 Constant batch size scheduler
$\text{[Constant BS] }b_{t}=b\text{ }(t\in\mathbb{N})$ (3)
does not always converge to 0 (a proof of Theorem 3.1 is given in Appendix A.3).
@@
Suppose that Assumption 1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that $\eta\leq\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L}$, and Constant BS defined by (3), where $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$ and $f^{\star}$ is the minimum value of $f$ over $\mathbb{R}^{d}$ (see (A1)). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{L\eta\sigma^{2}}{b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
@@
From the discussion in Section 2.2, we find that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2). The theoretical results for SHB can also be derived accordingly. Appendices A.2 and A.3 contain a detailed analysis of SHB.
@@ 3.2 Increasing batch size scheduler
An example of $b_{t}$ (Smith et al., 2018; Umeda and Iiduka, 2025) is, for all $m\in[0:M]$ and all $t\in S_{m}=\mathbb{N}\cap[\sum_{k=0}^{m-1}K_{k}E_{k},\sum_{k=0}^{m}K_{k}E_{k})$ ($S_{0}:=\mathbb{N}\cap[0,K_{0}E_{0})$),
$\text{[Exponential Growth BS] }b_{t}=\delta^{m\left\lceil\frac{t}{\sum_{k=0}^{m}K_{k}E_{k}}\right\rceil}b_{0},$ (4)
where $\delta>1$, and $E_{m}$ and $K_{m}$ are the numbers of, respectively, epochs and steps per epoch when the batch size is $\delta^{m}b_{0}$. For example, the exponential growth batch size defined by (4) with $\delta=2$ makes the batch size double every $E_{m}$ epochs. We may modify the parameters $a$ and $\delta$ to $a_{t}$ and $\delta_{t}$ monotone increasing with $t$. The total number of steps for the batch size to increase $M$ times is $T=\sum_{m=0}^{M}K_{m}E_{m}$.
@@
Suppose that Assumption 1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that
@@
and Exponential Growth BS as defined by (4) with $\delta>1$ and $\beta^{2}\delta>1$. Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{2L\eta\sigma^{2}K_{\max}E_{\max}\delta}{(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
@@
From the discussion in Section 2.2 indicating that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2), Theorem 3.2 leads to the following convergence rate of SHB defined by (2) with an increasing batch size.
Here, we sketch a proof of Theorem 3.2 (a detailed proof is given in Appendix A.2).
@@
    First, we show that $\frac{\sigma^{2}}{b_{t}}$ is an upper bound of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ (Proposition A.1) and that $(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ is an upper bound of the variance of $\bm{m}_{t}=(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f_{B_{i}}(\bm{\theta}_{i})$ (Lemma A.1) by using the idea underlying the proof of (Liu et al., 2020, Lemma 1).
@@
    Next, we show that an auxiliary point $\bm{z}_{t}=\frac{1}{1-\beta}\bm{\theta}_{t}-\frac{\beta}{1-\beta}\bm{\theta}_{t-1}$ ($t\geq 1$), which is used to analyze SGDM (Yan et al., 2018; Yu et al., 2019; Liu et al., 2020), satisfies $\mathbb{E}_{\bm{\xi}_{t}}[f(\bm{z}_{t+1})]\leq f(\bm{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\bm{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}$ by using the descent lemma (see (7)). Then, using the Cauchy–Schwarz inequality, Young’s inequality, and the upper bound of the variance of $\bm{m}_{t}$ (Lemma A.1), we arrive at an an upper bound on $\eta\mathbb{E}[X_{t}]$ (see (12)). As well, we find an upper bound on $\mathbb{E}[Y_{t}]$ by using the upper bound $\frac{\sigma^{2}}{b_{t}}$ of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ (see Lemma A.2 for details on the upper bounds of $\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$).
@@
    After that, we define the Lyapunov function $L_{t}$ by $L_{t}=f(\bm{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\|^{2}$, where $c_{i}$ is defined as in Lemma A.3. Using the above upper bounds of $\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$, we find that $\mathbb{E}[L_{t+1}-L_{t}]\leq-D\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+U_{t}$ (Lemma A.3), where $D\in\mathbb{R}$ depends on $\eta$, $\beta$, and $c_{1}$, and $U_{t}>0$ depends on $\sigma^{2}$, $b_{t}$, and $c_{1}$.
@@
    Then, setting $\eta$ such that it satisfies (5) (see also Appendix A.7) leads to the finding that $D\geq\frac{\eta}{2}>0$ and $U_{t}\leq L\eta^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ (see (19) and (20)). As a result, we have that
@@ 3.3 Setting of hyperparameter $\delta$ in Exponential Growth BS (4)
Let $\eta$ and $\beta$ be fixed in Algorithm 1 (e.g., $\eta=0.1$ and $\beta=0.9$). Then, Theorems 3.2 and A.2 indicate that $O(\sqrt{\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}+\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}})$ is an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for each of mini-batch NSHB and mini-batch SHB with exponential growth BS (4), which implies that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ becomes (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). In Section 4, we verify whether this theoretical result holds in practice.
3.4 Comparison of our convergence results with previous ones
@@
Let us compare Theorems 3.1, 3.2, A.1 and A.2 with the previous results listed in Table 1. Theorem 1 in Liu et al. (2020) ((6) in Table 1) indicated that NSHB using a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$. Since the upper bound $O(\sqrt{\frac{1}{T}+\sigma^{2}})$ converges to $O(\sigma)>0$ when $T\to+\infty$, NSHB in this case does not always converge to stationary points of $f$. The result in Liu et al. (2020) coincides with Theorem 3.1 indicating that NSHB has $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$ in the sense that NSHB using a constant learning rate and momentum weight does not converge to stationary points of ff111(Liu et al., 2020, Theorem 3) proved convergence of multistage SGDM. However, since the proof of (Liu et al., 2020, (60), Pages 35 and 36) might not hold for $\beta_{i}<1$, the theorem does not apply here.. Corollary 1 in Yu et al. (2019) ((5) in Table 1) indicated that SHB using constant $\eta=O(\frac{1}{\sqrt{T}})$ and constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/4}})$. $\eta=O(\frac{1}{\sqrt{T}})$ is needed in order set the number of steps $T$ before implementing SHB. Since $T$ is fixed, we cannot diverge it; that is, the upper bound $O(\frac{1}{T^{1/4}})$ for SHB is a fixed positive constant and does not converge to 0. Meanwhile, from Theorem 3.1, it follows that SHB with a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight also satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$. Hence, Theorem 3.1 coincides with the result in Corollary 1 in Yu et al. (2019) in the sense that SHB with a constant learning rate does not always converge to stationary points of $f$.
@@
Theorems 1 and 2 in Gitman et al. (2019) ((2) and (3) in Table 1) indicated that QHM, which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ or an increasing momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$. Our results in the form of Theorems 3.2 and A.2 guarantee the convergence of NSHB and SHB with constant learning rate $\eta=O(\frac{1}{L})$, constant momentum weight $\beta$, and an increasing batch size $b_{t}$ in the sense of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{\sqrt{T}})$.